OpenStudy (anonymous):

Factor the polynomial x^4 + 6x^2 - 7. Your answer can be written as (x+A)(x+B)(x^2+C) with integers A, B, C, and A < B. A = ? B = ? C = ?

OpenStudy (anonymous):

x^4 + 6x^2 - 7 =x^4 + 7x^2 - x^2- 7 =x^2(x^2+7) -1(x^2 + 7) =(x^2-1)(x^2+7) =(x-1)(x+1)(x^2+7) A<B A=-1 B=1 C=7